# Sample-efficient diffusion-based control of complex nonlinear systems

### v1

@@ Abstract
Complex nonlinear system control faces challenges in achieving sample-efficient, reliable performance. While diffusion-based methods have demonstrated advantages over classical and reinforcement learning approaches in long-term control performance, they are limited by sample efficiency. This paper presents SEDC (Sample-Efficient Diffusion-based Control), a novel diffusion-based control framework addressing three core challenges: high-dimensional state-action spaces, nonlinear system dynamics, and the gap between non-optimal training data and near-optimal control solutions. Through three innovations - Decoupled State Diffusion, Dual-Mode Decomposition, and Guided Self-finetuning - SEDC achieves 39.5%-49.4% better control accuracy than baselines while using only 10% of the training samples, as validated across three complex nonlinear dynamic systems. Our approach represents a significant advancement in sample-efficient control of complex nonlinear systems. The implementation of the code can be found at https://anonymous.4open.science/r/DIFOCON-C019.
@@ 1 Introduction
The control of complex systems plays a critical role across diverse domains, from industrial automation (Baggio et al., 2021) and biological networks (Gu et al., 2015) to robotics (Zhang et al., 2022). Given the challenges in deriving governing equations for empirical systems, data-driven control methods—which design control modules directly based on experimental data collected from the system, bypassing the need for explicit mathematical modeling—have gained prominence for their robust real-world applicability (Baggio et al., 2021; Janner et al., 2022; Ajay et al., 2022; Zhou et al., 2024; Liang et al., 2023; Wei et al., 2024; Ding et al., 2024).
Before data-driven control methods, traditional Proportional-Integral-Derivative (PID) (Li et al., 2006) controllers dominated complex system control through continuous error correction. However, these classical methods show limitations with complex nonlinear systems due to their linear control nature. Data-driven machine learning approaches have emerged to address these limitations by learning nonlinear control policies from interaction data, falling into three categories: supervised learning, reinforcement learning (RL), and diffusion-based methods. Supervised learning approaches like Behavior Cloning (BC) (Pomerleau, 1988) learn direct state-to-action mappings from expert demonstrations, while RL methods like Batch Proximal Policy Optimization (BPPO) (Zhuang et al., 2023) learn control policies through value function approximation, showing better adaptability to high-dimensional states than classical methods. However, both approaches often exhibit myopic decision-making in long-horizon tasks due to their iterative view of control dynamics. In contrast, diffusion-based methods (Janner et al., 2022; Ajay et al., 2022; Zhou et al., 2024; Liang et al., 2023; Wei et al., 2024) reformulate control as sequence generation, enabling comprehensive optimization over entire system trajectories. This long-term perspective allows diffusion-based methods to overcome limitations of both classical and RL approaches, achieving superior long-term control performance.
@@
The success of diffusion models in data-driven control stems from their exceptional ability to learn complex trajectory distributions from empirical data. In practice, these trajectories are typically collected from systems operated under empirical rules or random policies. Moreover, due to operational costs, the available data volume is often limited. Diffusion-based methods must therefore learn effective control policies from such non-optimal and sparse trajectory data—a challenge that manifests in three key aspects. First, limited data volume impedes sample-efficient learning in high-dimensional systems. Existing diffusion-based controllers (e.g., DiffPhyCon (Wei et al., 2024)) attempt to directly generate long-term ($T$ steps) state-action trajectories by learning a $T\times(P+M)$-dimensional distribution of system states $y^{P}$ and control inputs $u^{M}$. This joint distribution implicitly encodes system dynamics of state transitions under external control inputs, which often leads to physically inconsistent trajectories when training samples are insufficient. Second, learning control policies for nonlinear systems remains an open challenge both theoretically and practically. Traditional analytical methods (Baggio et al., 2021) designed for linear systems fail to perform robustly when applied to nonlinear systems. While diffusion-based approaches (Janner et al., 2022; Ajay et al., 2022; Zhou et al., 2024) employ deep neural networks (e.g., U-Net architectures) as denoising modules to capture nonlinearity, learning effective control policies from limited data remains particularly challenging for complex systems with strong nonlinearity, such as fluid dynamics and power grids. Third, extracting improved control policies from non-optimal training data poses fundamental difficulties. Diffusion-based methods (Janner et al., 2022) struggle when training data significantly deviates from optimal solutions. Although recent work (Wei et al., 2024) introduces reweighting mechanism to expand the solution space during generation, discovering truly near-optimal control policies remains elusive without explicit optimization guidance.
@@
To address these challenges, we propose SEDC (Sample-Efficient Diffusion-based Control), a novel diffusion-based framework for learning control policies of complex nonlinear systems with limited, non-optimal data. At its core, SEDC reformulates the control problem as a denoising diffusion process that samples control sequences optimized for reaching desired states while minimizing energy consumption. To address the curse of dimensionality, we introduce Decoupled State Diffusion (DSD), which confines diffusion process within state space and leverages inverse dynamics to generate control inputs, i.e., actions. This approach reduces learning complexity in high-dimensional systems while ensuring physics-aware control synthesis. To tackle strong nonlinearity, we propose Dual-Mode Decomposition (DMD) by designing a dual-UNet denoising module with residual connections. This architecture decomposes system dynamics into hierarchical linear and nonlinear components, enabling structured modeling of complex systems. To bridge the gap between non-optimal offline training data and optimal control policies, we introduce Guided Self-finetuning (GSF). This method progressively synthesizes guided control trajectories for iterative fine-tuning, facilitating exploration beyond initial training data and convergence toward near-optimal control strategies.
We demonstrate SEDC’s superiority over traditional, reinforcement learning, and diffusion-based methods through experiments on three typical complex nonlinear systems. Our model demonstrates 39.5%-49.4% improvement in control accuracy compared to state-of-the-art baselines while maintaining better balance between accuracy and energy consumption. In sample efficiency experiments, SEDC matches state-of-the-art performance using only 10% of the training samples. Additional ablation studies validate the effectiveness of SEDC’s key design components.
2 Related Works
Classic control methods. Data-driven control of complex systems has witnessed significant methodological developments across multiple paradigms. Classical control methods, represented by Proportional-Integral-Derivative (PID) controllers (Li et al., 2006), operate through continuous sensing-actuation cycles in a feedback-based manner. While these methods offer straightforward implementation, they face fundamental limitations when dealing with high-dimensional complex scenarios. More sophisticated analytical approaches, such as those presented in (Baggio et al., 2021), have attempted to determine optimal control inputs for complex networks without explicit dynamics knowledge. However, their foundation in linear systems theory inherently restricts their applicability to nonlinear systems.
Figure 1: Illustration of SEDC, the proposed conditional diffusion-based controller.
Data-driven control methods. The emergence of supervised learning (Pomerleau, 1988) and reinforcement learning (Haarnoja et al., 2018; Zhuang et al., 2023) has introduced more adaptive approaches to complex control problems, demonstrating promising results in sequential decision-making tasks. However, these approaches often struggle with real-world deployment due to computational constraints and the challenge of making effective decisions over extended time horizons. More recently, denoising diffusion probabilistic models (Ho et al., 2020) have emerged as a powerful framework for modeling high-dimensional distributions, achieving remarkable success across various domains including image, audio, and video generation (Dhariwal & Nichol, 2021; Kong et al., 2020; Ho et al., 2022). This success has inspired their application to control problems, with several works demonstrating their potential in robotic control (Janner et al., 2022; Ajay et al., 2022) and trajectory generation (Liang et al., 2023; Zhou et al., 2024). The diffusion framework has also shown promise in related technical domains such as optimization (Krishnamoorthy et al., 2023; Sun & Yang, 2023) and inverse problems (Chung et al., 2022). For diffusion-based control methods, works like (Janner et al., 2022; Ajay et al., 2022) demonstrate capabilities in generating long-term control trajectories for reinforcement learning environments, but they employ generic architectures that struggle to capture highly nonlinear dynamics, while our method incorporates specialized designs for effective nonlinear system learning. Other works like (Liang et al., 2023; Zhou et al., 2024) focus primarily on robotic control without specific considerations for complex system dynamics, while our approach is specifically designed to handle the challenges of high-dimensional state-action spaces and strong nonlinearity in complex systems. Recent work DiffPhyCon (Wei et al., 2024) incorporates reweighting techniques to optimize trajectories beyond the training data distribution, and attempts to optimize trajectories through implicit dynamics modeling and reweighting mechanisms, but this approach lacks explicit guidance for optimization and may lead to physically inconsistent predictions, whereas our method combines explicit dynamics modeling with guided optimization to ensure both physical consistency and optimality.
@@ 3.1 Problem Setting
[…]
$\mathbf{u}^{*}$ $=\arg\min_{\mathbf{u}}J(\mathbf{y},\mathbf{u})$ (1)
s.t. $\Psi(\mathbf{u},\mathbf{y})=0,\quad\mathbf{y}_{0}=\mathbf{y}_{0}^{*},\quad\mathbf{y}_{T}=\mathbf{y}_{f},$
where $\mathbf{y}\in\mathbb{R}^{T\times N}$ is the corresponding complete state trajectory given $\mathbf{y}_{0}$ and $\Psi(\mathbf{u},\mathbf{y})=0$. Here, $\Psi(\mathbf{u},\mathbf{y})=0$ represents the system dynamics constraint implicitly defined by dataset $D$. This constraint effectively serves as a data-driven representation of the unknown dynamics equation $\dot{\mathbf{y}}_{t}=\Phi(\mathbf{y}_{t},\mathbf{u}_{t})$.
@@ 3.2 Diffusion Model
Diffusion models have become leading generative models, showing exceptional results across image synthesis, audio generation and other applications (Ho et al., 2020; Dhariwal & Nichol, 2021; Song & Ermon, 2019). These models, when applied to trajectory generation, operate by progressively adding noise to sequential data in the forward process and then learning to reverse this noise corruption through a denoising process. We denote that $\mathbf{x}^{k}$ represents the sequential data at diffusion timestep $k$. In the forward process, a clean trajectory $\mathbf{x}^{0}$ is progressively corrupted through $K$ timesteps, resulting in a sequence of increasingly noisy versions $\mathbf{x}^{1},\mathbf{x}^{2},...,\mathbf{x}^{K}$. Each step applies a small amount of Gaussian noise:
$q(\mathbf{x}^{k}|\mathbf{x}^{k-1})=\mathcal{N}(\mathbf{x}^{k};\sqrt{1-\beta^{k}}\mathbf{x}^{k-1},\beta^{k}\mathbf{I}),$
where $\beta^{k}$ is a variance schedule that controls the noise level. With a large enough $K$ we can get $q(\mathbf{x}^{K})\approx\mathcal{N}(\mathbf{x}^{K};\mathbf{0},\mathbf{I})$. In the reverse process, the diffusion model learns to gradually denoise the data, starting from pure noise $\mathbf{x}^{K}$ and working backward to reconstruct the original plausible trajectory $\mathbf{x}^{0}$. Each denoising step is conditioned on the start and target state:
$p_{\theta}(\mathbf{x}^{k-1}|\mathbf{x}^{k},\mathbf{y}_{0}^{*},\mathbf{y}_{f})=\mathcal{N}(\mathbf{x}^{k-1};\mathbf{\mu}_{\theta}(\mathbf{x}^{k},k,\mathbf{y}_{0}^{*},\mathbf{y}_{f}),\mathbf{\Sigma}^{k}),$
where $\theta$ represents the learnable parameters of the model and $\Sigma^{k}$ is from a fixed schedule.
Training of diffusion model. In order to facilitate the design of the denoising network, the network with $\theta$ for the denoising process does not directly predict $\mathbf{\mu}$. Instead, it is trained to learn to predict clean trajectory $\mathbf{x}^{0}$, outputting $\hat{\mathbf{x}}^{0}$.
The training objective for diffusion models typically involves minimizing the variational lower bound (VLB) on the negative log-likelihood (Sohl-Dickstein et al., 2015). In practice, this often reduces to a form of denoising score matching (Song & Ermon, 2019):
$\mathbb{E}_{\mathbf{x},k,\mathbf{y}_{0}^{*},\mathbf{y}_{f},\mathbf{\epsilon}}[||\boldsymbol{\mathbf{x}}-\boldsymbol{\mathbf{x}}_{\theta}(\boldsymbol{\mathbf{x}}^{k},k,\mathbf{y}_{0}^{*},\mathbf{y}_{f})||^{2}],$
where $\mathbf{x},k,\mathbf{y}_{0}^{*},\mathbf{y}_{f}$ are sampled from the dataset, $k\sim\mathcal{U}\{1,2,...,K\}$ is the step index and $\mathbf{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is the noise used to corrupt $\mathbf{x}$.

### v2

# Sample-Efficient Diffusion-based Control of Complex Physics Systems

@@ Abstract
Controlling complex physics systems is important in diverse domains. While diffusion-based methods have demonstrated advantages over classical model-based approaches and myopic sequential learning methods in achieving global trajectory consistency, they are limited by sample efficiency. This paper presents SEDC (Sample-Efficient Diffusion-based Control), a novel framework addressing core challenges in complex physics systems: high-dimensional state-control spaces, strong nonlinearities, and the gap between non-optimal training data and near-optimal control laws. Our approach introduces a novel control paradigm by architecturally decoupling state-control modeling and decomposing dynamics, while a guided self-finetuning process iteratively refines the control law towards optimality. We validate SEDC across diverse complex nonlinear systems, including high-dimensional fluid dynamics (Burgers), chaotic synchronization networks (Kuramoto), and real-world power grid stability control (Swing Equation). Our method achieves 39.5%-47.3% better control accuracy than state-of-the-art baselines while using only 10% of the training samples. The implementation is available at here.
@@ 1 Introduction
The control of complex physics systems plays a critical role across diverse domains, from industrial automation (Baggio et al., 2021) to biological networks (Gu et al., 2015). Given the challenges in deriving governing equations for empirical systems, data-driven control methods—which design control modules directly based on experimental data collected from the system, bypassing the need for explicit mathematical modeling—have gained prominence for their robust real-world applicability (Baggio et al., 2021; Janner et al., 2022; Ajay et al., 2022; Zhou et al., 2024; Liang et al., 2023; Wei et al., 2024b).
Traditional Proportional-Integral-Derivative (PID) (Li et al., 2006) and Model Predictive Control (MPC) (Schwenzer et al., 2021) methods are limited in complex nonlinear systems. PID controllers struggle with nonlinearities, while MPC’s performance depends on model accuracy and is computationally intensive for long-horizon tasks. Data-driven approaches have emerged to address these issues, including supervised learning, sequential learning approaches, and diffusion-based methods. Supervised trajectory fitting baselines (e.g., BC (Pomerleau, 1988)) and sequential learning methods (e.g., BPPO (Zhuang et al., 2023)) often make myopic, step-by-step decisions, leading to suboptimal outcomes in long-horizon tasks. In contrast, diffusion-based methods (Janner et al., 2022; Ajay et al., 2022; Zhou et al., 2024; Liang et al., 2023; Wei et al., 2024b) treat control as a global trajectory generation problem. By generating the entire control plan in a single sample, they achieve comprehensive optimization over the full trajectory, avoiding the pitfalls of iterative methods and enabling superior long-term performance.
@@
The success of diffusion models is dependent on their ability to learn complex trajectory distributions. However, in practice, the available trajectory data is often non-optimal and sparse due to collection under empirical rules or random actuation and high operational costs. This presents a key challenge for diffusion-based methods: synthesizing effective control laws from limited and suboptimal data. First, limited data volume impedes sample-efficient modeling in high-dimensional physics systems. Existing diffusion-based controllers (Wei et al., 2024a, b; Hu et al., 2025) attempt to directly generate long-term ($T$ steps) state-control trajectories by learning a $T\times(P+M)$-dimensional distribution of system states $y^{P}$ and control inputs $u^{M}$. This joint distribution implicitly encodes system dynamics of state transitions under external control inputs, which often leads to physically inconsistent trajectories when training samples are insufficient (Janner et al., 2022). Second, synthesizing control laws for nonlinear systems remains an open challenge both theoretically and practically. Traditional analytical methods (Baggio et al., 2021) designed for linear systems fail to perform robustly when applied to nonlinear systems. While diffusion-based approaches (Janner et al., 2022; Ajay et al., 2022; Zhou et al., 2024; Zhong et al., 2025; Hu et al., 2025) employ deep neural networks (e.g., U-Net architectures) as denoising modules to capture nonlinearity, synthesizing effective control laws from limited data remains particularly challenging for complex systems with strong nonlinearity, such as fluid dynamics (Brunton et al., 2020) and power grids (Baggio et al., 2021). Third, extracting improved control laws from non-optimal training data poses fundamental difficulties. Diffusion-based methods (Janner et al., 2022) struggle when training data significantly deviates from optimal solutions. Although recent work (Wei et al., 2024b) introduces reweighting mechanism to expand the solution space during generation, discovering truly near-optimal control laws remains elusive without explicit optimization guidance.
@@
To address these challenges, we propose SEDC (Sample-Efficient Diffusion-based Control), a novel diffusion-based framework for synthesizing control laws of complex physics systems with limited, non-optimal data. At its core, SEDC reformulates the control problem as a denoising diffusion process that samples control sequences optimized for reaching desired states while minimizing energy consumption. We then solve the sample efficiency challenge by addressing its three key aspects. To address the curse of dimensionality, we introduce Decoupled State Diffusion (DSD), which simplifies the modeling complexity of the generative task. By diffusing only on the more structured state space, rather than the complex joint state-control space, DSD achieves higher sample efficiency. A separate inverse dynamics model is then used to ensure physical consistency. To tackle strong nonlinearity, we propose Dual-Mode Decomposition (DMD) by designing a dual-UNet denoising module with residual connections. This architecture decomposes system dynamics into hierarchical linear and nonlinear components, enabling structured modeling of complex systems. To bridge the gap between non-optimal training data and optimal control laws, we introduce the Guided Self-finetuning (GSF) mechanism, which progressively synthesizes guided control trajectories for iterative finetuning, facilitating manifold expansion beyond initial training data and convergence toward near-optimal control laws.
Our contributions are summarized as follows:
  • We introduce a data-driven framework that significantly enhances sample efficiency in controlling high-dimensional nonlinear systems via diffusion models.
  • Experiments on three complex nonlinear systems show that SEDC outperforms traditional, sequential, and diffusion-based baselines, achieving a 39.5%–47.3% improvement in accuracy with a better accuracy-energy balance.
  • SEDC matches top-tier performance using only 10% of training data. We further demonstrate its scalability on 2D PDEs and robustness to non-invertible systems.
  • We confirm the method’s computational efficiency and provide extensive ablation studies to verify the specific contribution of each architectural component.
Figure 1: Overview of SEDC. The framework consists of a training/finetuning (top panel), inference process (middle panel) and finetuning process (bottom panel). The core denoising network employs our Dual-Mode Decomposition (DMD) architecture (right panel). Both training and inference leverage Decoupled State Diffusion (DSD) by diffusing only on states and using a separate inverse dynamics model to recover controls.
2 Related Work
Data-driven control encompasses various paradigms, which can be broadly categorized by their approach to trajectory generation. One major paradigm is iterative, feedback-based control. Classical methods like PID controllers (Li et al., 2006) operate via real-time error correction but face limitations in high-dimensional complex scenarios. More contemporary approaches center on system identification, such as Dynamic Mode Decomposition (Tu, 2013) and Koopman operator theory (Mauroy et al., 2020). These methods first learn an explicit dynamics model from data and then design a controller, often within a Model Predictive Control (MPC) framework (Schwenzer et al., 2021). While robust for real-time adaptation, this two-stage paradigm is susceptible to compounding errors, particularly in sample-scarce settings. Inaccuracies in the learned model can accumulate over long horizons, degrading control performance. Sequential learning approaches such as supervised trajectory fitting (Pomerleau, 1988) and reinforcement learning (Haarnoja et al., 2018; Zhuang et al., 2023) offer adaptive methods but can struggle with long-horizon credit assignment and compounding errors due to their myopic, step-by-step nature (Ross et al., 2011).
In contrast, global trajectory planning reframes control as a holistic generation problem, for which denoising diffusion models (Ho et al., 2020; Dhariwal and Nichol, 2021; Kong et al., 2020; Ho et al., 2022) have emerged as a powerful tool. By generating the entire control plan as a single, coherent sample, these methods capture long-term dependencies and avoid the pitfalls of iterative error accumulation. Seminal works (Janner et al., 2022; Ajay et al., 2022) demonstrated this potential in robotics, but their generic architectures struggle with strong nonlinearities; our Dual-Mode Decomposition (DMD) architecture addresses this with a structured inductive bias. Subsequent research has tackled specific limitations. DiffPhyCon (Wei et al., 2024b) uses reweighting to synthesize trajectories extending beyond the training distribution. This approach, however, requires training separate denoising networks to model decomposed energy functions (one for the prior and one for the conditional distribution). Moreover, its joint state-control modeling can exacerbate the curse of dimensionality, which our Decoupled State Diffusion (DSD) alleviates by diffusing over the state space alone. To bridge the data-optimality gap, AdaptDiffuser (Liang et al., 2023) fine-tunes on discriminator-filtered trajectories, whereas our Guided Self-finetuning (GSF) employs a simpler, filter-free loop. Beyond standard architectures, advanced approaches have emerged to address specific physical challenges—such as WDNO (Hu et al., 2024) for multi-resolution dynamics and SafeDiffCon (Hu et al., 2025) for safety constraints. However, these methods typically prioritize simulation fidelity or boundary satisfaction, rather than the sample-efficient synthesis of near-optimal control policies from sparse, suboptimal data.
@@ 3.1 Problem Setting
The dynamics of a controlled complex system can be represented by the differential equation $\dot{\mathbf{y}}_{t}=\Phi(\mathbf{y}_{t},\mathbf{u}_{t})$, where $\mathbf{y}_{t}\in\mathbb{R}^{N}$ represents the system state and $\mathbf{u}_{t}\in\mathbb{R}^{M}$ denotes the control input. We assume the system satisfies the controllability condition without loss of generality: for any initial state $\mathbf{y}_{0}^{*}$ and target state $\mathbf{y}_{f}$, there exists a finite time $T$ and a corresponding control input $\mathbf{u}$ that can drive the system from $\mathbf{y}_{0}^{*}$ to $\mathbf{y}_{f}$. This assumption ensures the technical feasibility of our control objectives. In practical applications, beyond achieving state transitions, we need to optimize the energy consumption during the control process. The energy cost can be quantified using the L2-norm integral of the control input: $J(\mathbf{y},\mathbf{u})=\int_{0}^{T}|\mathbf{u}(t^{\prime})|^{2}dt^{\prime}$. Consider a dataset $D=\{\mathbf{u}^{(i)},\mathbf{y}^{(i)}\}_{i=1}^{P}$ containing $P$ non-optimal control trajectories, where each trajectory consists of: (1) complete state trajectories $\mathbf{y}^{(i)}={\mathbf{y}^{(i)}_{0},...,\mathbf{y}^{(i)}_{T}}$ sampled at fixed time intervals; (2) corresponding control input sequences $\mathbf{u}^{(i)}={\mathbf{u}^{(i)}_{0},...,\mathbf{u}^{(i)}_{T-1}}$. Our objective is satisfying the boundary constraint (reaching $\mathbf{y}_{f}$ from $\mathbf{y}_{0}^{*}$) and minimize the energy cost $J(\mathbf{y},\mathbf{u})$ within the feasible solution space. Formally, this is a constrained optimization problem:
$\min_{\mathbf{u}}$ $J(\mathbf{y},\mathbf{u})$ (1)
s.t. $\mathbf{y}_{T}=\mathbf{y}_{f}\quad\text{(Target Satisfaction)},$
$\Psi(\mathbf{u},\mathbf{y})=0\quad\text{(Dynamics Consistency)},$
where $\mathbf{y}\in\mathbb{R}^{T\times N}$ is the corresponding complete state trajectory given $\mathbf{y}_{0}=\mathbf{y}_{0}^{*}$. $\Psi(\mathbf{u},\mathbf{y})=0$ represents the system dynamics constraint implicitly defined by dataset $D$. This constraint effectively serves as a data-driven representation of the unknown dynamics equation $\dot{\mathbf{y}}_{t}=\Phi(\mathbf{y}_{t},\mathbf{u}_{t})$.
@@ 3.2 Diffusion Model
[…]